alesuarez:

A bus is traveling with a uniform acceleration of 2.75 meters/second2. If the initial velocity of the bus is 16.5 meters/second, what will its velocity be after 9.00 seconds?

justjm:

Use your kinematics equations \(v_f=v_i-at\) \(a=2.75 m/s^2\) \(v_i = 16.5m/s\) \(t=9.00 s\) \(v_f=v_f\) Plug in and solve for \(v_f\)